OpenStudy (prominer10000):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A. 2 B. -4/5 C. 5/4 D. -2 I'm not sure what this would be.. Please help!

OpenStudy (jiteshmeghwal9):

\[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (jiteshmeghwal9):

Y2=1. X2=4 X1=1 y1=-5

OpenStudy (prominer10000):

So would it be B?

OpenStudy (prominer10000):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

\[\frac{1-(-5)}{4-1}=?\]

OpenStudy (prominer10000):

6/3 right?

OpenStudy (jiteshmeghwal9):

Right

OpenStudy (jiteshmeghwal9):

Got it ?

OpenStudy (prominer10000):

So what wold the answer be though? Is 6/3 the slope or what?

OpenStudy (jiteshmeghwal9):

It is the slope

OpenStudy (jiteshmeghwal9):

6/3=2=slope

OpenStudy (triciaal):

always reduce to the lowest term when you can

OpenStudy (prominer10000):

Ah ok so it would be A.

OpenStudy (triciaal):

yes